# Compositional Generative Multiphysics and Multi-component Simulation

### v1

@@ Appendix B Additional details for reaction-diffusion
Model structure. The 2D U-Net and 2D FNO serve as both the surrogate and MultiSimDiff. U-Net consists of modules: a downsampling encoder, a middle module, and an upsampling decoder. The encoder and decoder comprise four layers, each with three residual modules and downsampling/upsampling convolutions, with the third module incorporating attention mechanisms. The middle module also contains three residual modules, with attention mechanisms included in the second module. The input data is encoded into a hidden dimension before undergoing sequential downsampling and upsampling. FNO consists of three modules: a lift-up encoder, $n$ FNO layers, and a projector decoder. Each FNO layer includes a spectral convolution, a spatial convolution, and a layer normalization. The surrogate model predicts the evolution of the current physical process using its initial conditions and those of other physical processes. Its input dimension is [b, 1, 10, 20] and output dimension is [b, 1, 10, 20]. The diffusion model has an input dimension of [b, 2, 10, 20] and an output dimension of [b, 1, 10, 20], with b representing the batch size. The shape of initial condition of [b, 1, 1, 20] and will repeat to align the required shape. The diffusion step of the diffusion model is set to 250. More details are shown in Table 4.
Training. The surrogate model and MultiSimDiff are trained similarly, with further details in Table 5.
Inference. The hyperparameter $K$ is set to 2. The surrogate models’ combination algorithm in experiments 1 and 2 is identical, as demonstrated in Algorithm 3. The relaxation factor $\alpha$ is set to 0.5.
Algorithm 3 Surrogate model combination algorithm.
@@
  Initialize constant fields $z_{i}$, $m=0$
  while $m<M$ and $\epsilon>\epsilon_{max}$ do
     $m=m+1$
     $\hat{z}_{i}=z_{i}$
     for $i=1,...,N$ do
        $z_{i}=\alpha\epsilon_{\theta}^{i}(z_{\neq i},C)+(1-\alpha)\hat{z}_{i}$
     end for
     $\epsilon=L_{1}(z_{i}-\hat{z}_{i})$
  end while
  return  ${z_{i}}$
@@ Appendix C Additional details for nuclear thermal coupling
Model structure. The 3D U-Net and 3D FNO serve as both the surrogate model and MultiSimDiff, using a layer design identical to the 2D. For regional coupling, concatenation is directly applied to the channel dimension using the concat function. In contrast, for interface coupling, dimensions must be replicated to align spatially before concatenation. The conditioning of the diffusion step for FNO is operating in spectrum space (Gupta & Brandstetter, 2023), which is better than in the original space for this problem. The diffusion step of the diffusion model is set to 250. More details are shown in Table 6.
Training. The surrogate model and MultiSimDiff are trained similarly, but training neutron physics fields using MultiSimDiff requires more time to converge, with further details in Table 7.
@@
Detailed results. Fig. 5 presents the results of predicting various physical fields using the last time step surrogate model and MultiSimDiff + U-Net on the final test data. The neutron physics field and solid temperature field are represented by $\phi$ and $T_{s}$, respectively. The fluid fields include four physical quantities: $T_{f},P,u_{x},u_{y}$, totaling six quantities. Since the neutron physics field and the $u_{y}$ component of the fluid fields are less influenced by other physical processes, the surrogate model can still make predictions, but the accuracy is lower than that of the MultiSimDiff. Besides, the surrogate model has failed to predict the other physical processes. In contrast, MultiSimDiff continues to provide relatively accurate predictions, although some distortions are observed in certain regions.
@@
Figure 5: Comparison of surrogate model and MultiSimDiff for predicting all physical fields.
@@ Appendix D Additional details for prismatic fuel element
Problem description. This problem aims to predict the thermal and mechanical performance of prismatic fuel elements in heat pipe reactor Ma et al. (2022) at different source power. The reactor core is stacked up using a hexagonal prism SiC matrix, with multiple holes dispersed in the matrix for containing fuel elements and heat pipes as shown in Fig. 6. The SiC matrix plays a role in locating the fuel and heat pipes at expected positions in the core. The entire structure consists of two basic components, one oriented upwards and the other downwards, as illustrated in Fig. 7. Fission energy released in fuel elements is dissipated using heat pipes. Both the fuel elements and heat pipes are considered as boundaries here, and only the more concerned matrix behavior is analyzed in the demonstration. Only strain is predicted here since stress can be derived from the mechanical constitutive equation, and displacement is obtained through strain integration. The analysis uses the plane strain assumption ($\varepsilon_{z}=0$) and excludes irradiation effects, simplifying it to a steady-state problem.
@@
Model structure. The Geo-FNO and Transolver serve as both the surrogate model and MultiSimDiff. Geo-FNO enhances FNO for irregular meshes using three modules: a geometry encoder that converts physical fields from irregular to latent uniform meshes, FNO functioning in latent space, and a geometry decoder that transforms physical fields from the uniform mesh back to the original irregular mesh. We utilize a 2D Geo-FNO that transforms into a 2D uniform mesh. Transolver is designed to tackle complex structural simulation problems involving numerous mesh points by learning the intrinsic physical states of the discretized domain. Given a mesh set with $N$ points and $C$ features per point, the network first assigns each mesh point to $M$ potential slices, transforming the shape from $N\times C$ to $M\times N\times C$. It then applies spatially weighted aggregation, resulting in a shape of $M\times C$. Self-attention is used to capture intricate correlations among different slices, after which the data is transformed back to the mesh points. The conditioning of diffusion step for Geo-FNO is also operating in spectrum space (Gupta & Brandstetter, 2023). More details about the network can be found in (Wu et al., 2024a). The setting of hyperparameters is shown in Table 8. The diffusion step of the diffusion model is set to 250.
@@
Training. The surrogate model and MultiSimDiff are trained similarly, but training MultiSimDiff requires more time to converge, with further details in Table 9.
@@
Inference. This problem uses the same neural network to predict the performance of all elements, allowing for simultaneous updates of the physical fields and enhancing inference speed. This method applies to both diffusion and surrogate models. The hyperparameter $K$ is set to 3.
@@
Detailed results. Fig. 8 and Fig. 9 compares the results of predicting the large structure using the surrogate model and MultiSimDiff + U-Net. Because the surrogate model of FNO fails in predicting large structures, only the results of MultiSimDiff + FNO are provided in Fig. 10. The strain is only displayed in the $x$-direction due to its similarity in both $x$ and $y$. The error graph indicates that MultiSimDiff offers more accurate predictions.
@@
Graph neural network and Graph Transformer configuration. For GIN and SAN, each component is treated as a node in the graph, with training conducted on a small graph of 16-component, and ultimately tested on a larger graph of 64-component. Compared with the surrogate model and MultiSimDiff, they use only the system’s input as input features. In contrast, the surrogate model and MultiSimDiff enrich its input by incorporating the solutions from the surrounding component, thereby improving accuracy, as demonstrated in Table 3. The input to the GIN and SAN is the heat flux density and boundary conditions of each component, and the output is the physical quantities at all grid points on the component. GIN updates the nodes on the graph through the graph structure, whereas SAN captures graph structural information by inputting the eigenvalues and eigenvectors of the graph Laplacian matrix into a transformer. The training settings of GIN and SAN are consistent with the MultiSimDiff. We have adjusted the number of network layers and the size of hidden layers to obtain the model with optimal performance.
@@
(b) MultiSimDiff when predicting $T$.
Figure 8: Comparison of surrogate models and MultiSimDiff + U-Net for predicting the temperature of large structures.
@@
(b) MultiSimDiff when predicting $\varepsilon_{x}$.
Figure 9: Comparison of surrogate models and MultiSimDiff + U-Net for predicting the strain of large structures.
(a) MultiSimDiff when predicting $T$.
(b) MultiSimDiff when predicting $\varepsilon_{x}$.
Figure 10: The results of MultiSimDiff + FNO for predicting large structure.
@@ F.3 Selection of Hyperparameter of $\lambda$
This section compares the computational efficiency of MultiSimDiff, surrogate model, and numerical programs. The time unit for each experiment is defined as the time required for a single neural network inference. These three experiments all use the most accurate model. Since the surrogate model and MultiSimDiff both use the same network architecture and have consistent network parameters, it is assumed that the time for a single inference using these two methods is equal. The numerical programs are run on the CPU and have all been optimized to the best parallel count.
@@
The results are presented in Table 20. In experiment 1, the problem is relatively simple, and the numerical algorithm achieves efficient solutions through explicit time stepping, while the introduction of MultiSimDiff actually reduces efficiency. However, in experiment 2, which addresses more complex problems, MultiSimDiff achieves a 29-fold acceleration compared to numerical programs. In experiment 3, comparing the results of 16 components with 64 components, it is observed that as the computational scale increases, the acceleration effect of MultiSimDiff becomes increasingly significant. Furthermore, when dealing with multi-component problems, the surrogate model requires iteration to ensure the convergence of solutions across all components. Due to the large number of components, the number of iterations needed significantly increases compared to multiphysics problems, resulting in higher efficiency for MultiSimDiff. In addition, we have only compared the efficiency of single computations for all experiments. When dealing with multiple problems simultaneously, the acceleration provided by MultiSimDiff will be even more pronounced due to the parallel nature of GPU computing.
In general, the more complex the problem, the more pronounced the acceleration effect of MultiSimDiff becomes. In fact, the problems in experiment 2 and experiment 3 have been simplified to a certain extent, and the actual situations are even more complex. Therefore, MultiSimDiff holds significant value in solving real-world complex engineering problems.
@@
Experiment Unit (s) Numerical program Surrogate model MultiSimDiff Speedup
@@
In this section, we discuss the application scenarios of MultiSimDiff for multi-component simulation from both theoretical and practical perspectives.
From a theoretical perspective, in the derivation of Section 3.2, we make an assumption: the solution on a multi-component structure is an undirected graph that satisfies the local Markov property, meaning that any two non-adjacent variables are conditionally independent given all other variables. Using this property, we derived Eq. 14. We believe this assumption is applicable to most problems because physical fields are continuous in space, and the information exchange between any two points must be transmitted through the points in between. However, there is a class of problems to which current methods cannot be directly applied, which is the partial differential equation that requires determining eigenvalues:

### v2

@@ Appendix B Additional details for reaction-diffusion
Model structure. The 2D U-Net and 2D FNO serve as both the surrogate and M2PDE. U-Net consists of modules: a downsampling encoder, a middle module, and an upsampling decoder. The encoder and decoder comprise four layers, each with three residual modules and downsampling/upsampling convolutions, with the third module incorporating attention mechanisms. The middle module also contains three residual modules, with attention mechanisms included in the second module. The input data is encoded into a hidden dimension before undergoing sequential downsampling and upsampling. FNO consists of three modules: a lift-up encoder, $n$ FNO layers, and a projector decoder. Each FNO layer includes a spectral convolution, a spatial convolution, and a layer normalization. The surrogate model predicts the evolution of the current physical process using its initial conditions and those of other physical processes. Its input dimension is [b, 1, 10, 20] and output dimension is [b, 1, 10, 20]. The diffusion model has an input dimension of [b, 2, 10, 20] and an output dimension of [b, 1, 10, 20], with b representing the batch size. The shape of initial condition of [b, 1, 1, 20] and will repeat to align the required shape. The diffusion step of the diffusion model is set to 250. More details are shown in Table 4.
Training. The surrogate model and M2PDE are trained similarly, with further details in Table 5.
Inference. The hyperparameter $K$ is set to 2. The surrogate models’ combination algorithm in experiments 1 and 2 is identical, as demonstrated in Alg 3. The relaxation factor $\alpha$ is set to 0.5.
Algorithm 3 Surrogate model combination algorithm for multiphysics simulation.
@@
1:  Initialize constant fields $z_{i},i=1,...,N$, $m=0$
2:  while $m<M$ and $\epsilon>\epsilon_{max}$ do
3:     $m=m+1$
4:     for $i=1,...,N$ do
5:        $\hat{z}_{i}=z_{i}$
6:        $z_{i}=\alpha\epsilon_{\theta}^{i}(z_{\neq i},C)+(1-\alpha)\hat{z}_{i}$
7:     end for
8:     $\epsilon=L_{1}(z_{i}-\hat{z}_{i}),i=1,...,N$
9:  end while
9:  $z_{1},z_{2},...,z_{N}$
@@ Appendix C Additional details for nuclear thermal coupling
Model structure. The 3D U-Net and 3D FNO serve as both the surrogate model and M2PDE, using a layer design identical to the 2D. For regional coupling, concatenation is directly applied to the channel dimension using the concat function. In contrast, for interface coupling, dimensions must be replicated to align spatially before concatenation. The conditioning of the diffusion step for FNO is operating in spectrum space (Gupta & Brandstetter, 2023), which is better than in the original space for this problem. The diffusion step of the diffusion model is set to 250. More details are shown in Table 6.
Training. The surrogate model and M2PDE are trained similarly, but training neutron physics fields using M2PDE requires more time to converge, with further details in Table 7.
@@
Detailed results. Fig. 5 llustrates how the relative errors are distributed along the $x,y,t$ direction. Overall, the neutron field exhibits a relatively uniform distribution of errors along all three directions. In $x$ direction, the error in both the solid and the fluid slightly increases near their interface, likely due to coupling effects in this region. Along the $y$ direction, the solid temperature error gradually increases with distance, whereas the fluid error shows the opposite trend, which may be related to the distinct thermal and physical properties of the two media. Fig. 6 presents the results of predicting various physical fields using the last time step surrogate model and M2PDE + U-Net on the final test data. The neutron physics field and solid temperature field are represented by $\phi$ and $T_{s}$, respectively. The fluid fields include four physical quantities: $T_{f},P,u_{x},u_{y}$, totaling six quantities. Since the neutron physics field and the $u_{y}$ component of the fluid fields are less influenced by other physical processes, the surrogate model can still make predictions, but the accuracy is lower than that of the M2PDE. Besides, the surrogate model has failed to predict the other physical processes. In contrast, M2PDE continues to provide relatively accurate predictions, although some distortions are observed in certain regions.
@@
(a) t direction.
(b) x direction.
(c) y direction.
Figure 5: The distribution of errors along the x, y, and t directions.
@@
Figure 6: Comparison of surrogate model and M2PDE for predicting all physical fields.
@@ Appendix D Additional details for prismatic fuel element
Problem description. This problem aims to predict the thermal and mechanical performance of prismatic fuel elements in heat pipe reactor (Ma et al., 2022) at different source power. The reactor core is stacked up using a hexagonal prism SiC matrix, with multiple holes dispersed in the matrix for containing fuel elements and heat pipes as shown in Fig. 7. The SiC matrix plays a role in locating the fuel and heat pipes at expected positions in the core. The entire structure consists of two basic components, one oriented upwards and the other downwards, as illustrated in Fig. 8. Fission energy released in fuel elements is dissipated using heat pipes. Both the fuel elements and heat pipes are considered as boundaries here, and only the more concerned matrix behavior is analyzed in the demonstration. Only strain is predicted here since stress can be derived from the mechanical constitutive equation, and displacement is obtained through strain integration. The analysis uses the plane strain assumption ($\varepsilon_{z}=0$) and excludes irradiation effects, simplifying it to a steady-state problem.
@@
Model structure. The Geo-FNO and Transolver serve as both the surrogate model and M2PDE. Geo-FNO enhances FNO for irregular meshes using three modules: a geometry encoder that converts physical fields from irregular to latent uniform meshes, FNO functioning in latent space, and a geometry decoder that transforms physical fields from the uniform mesh back to the original irregular mesh. We utilize a 2D Geo-FNO that transforms into a 2D uniform mesh. Transolver is designed to tackle complex structural simulation problems involving numerous mesh points by learning the intrinsic physical states of the discretized domain. Given a mesh set with $N$ points and $C$ features per point, the network first assigns each mesh point to $M$ potential slices, transforming the shape from $N\times C$ to $M\times N\times C$. It then applies spatially weighted aggregation, resulting in a shape of $M\times C$. Self-attention is used to capture intricate correlations among different slices, after which the data is transformed back to the mesh points. The conditioning of diffusion step for Geo-FNO is also operating in spectrum space (Gupta & Brandstetter, 2023). More details about the network can be found in (Wu et al., 2024a). The setting of hyperparameters is shown in Table 8. The diffusion step of the diffusion model is set to 250.
@@
Training. The surrogate model and M2PDE are trained similarly, but training M2PDE requires more time to converge, with further details in Table 9.
@@
Inference. This problem uses the same neural network to predict the performance of all elements, allowing for simultaneous updates of the physical fields and enhancing inference speed. This method applies to both diffusion and surrogate models. The surrogate models’ combination algorithm for mult-component simulation is demonstrated in Alg 4, the relaxation factor $\alpha$ is set to 0.5. The hyperparameter $K$ is set to 3.
@@
Detailed results. Fig. 9 and Fig. 10 compares the results of predicting the large structure using the surrogate model and M2PDE + Transolver. Because the surrogate model of FNO fails in predicting large structures, only the results of M2PDE + FNO are provided in Fig. 11. The strain is only displayed in the $x$-direction due to its similarity in both $x$ and $y$. The error graph indicates that M2PDE offers more accurate predictions.
@@
Graph neural network and Graph Transformer configuration. For GIN, SAN and MeshGraphNet, each component is treated as a node in the graph, with training conducted on a small graph of 16-component, and ultimately tested on a larger graph of 64-component. Compared with the surrogate model and M2PDE, they use only the system’s input as input features. In contrast, the surrogate model and M2PDE enrich its input by incorporating the solutions from the surrounding component, thereby improving accuracy, as demonstrated in Table 3. The input to the GIN, SAN and MeshGraphNet is the heat flux density and boundary conditions of each component, and the output is the physical quantities at all grid points on the component. GIN updates the nodes on the graph through the graph structure, whereas SAN captures graph structural information by inputting the eigenvalues and eigenvectors of the graph Laplacian matrix into a transformer. The training settings of GIN, SAN and MeshGraphNet are consistent with the M2PDE. We have adjusted the number of network layers and the size of hidden layers to obtain the model with optimal performance.
Algorithm 4 Surrogate model combination algorithm for multi-component simulation.
0:  A surrogate model $\epsilon_{\theta}(z_{\partial v},C)$, outer inputs $C$, maximum number of iterations $M$, tolerance $\epsilon_{max}$, relaxation factor $\alpha$.
1:  Initialize constant fields $z_{v_{i}},i=1,...,N$, $m=0$
2:  while $m<M$ and $\epsilon>\epsilon_{max}$ do
3:     $m=m+1$
4:     for $i=1,...,N$ do
5:        $\hat{z}_{v_{i}}=z_{v_{i}}$
6:        $z_{v_{i}}=\alpha\epsilon_{\theta}(z_{\partial v_{i}},C)+(1-\alpha)\hat{z}_{v_{i}}$
7:     end for
8:     $\epsilon=L_{1}(z_{\partial v_{i}}-\hat{z}_{\partial v_{i}}),i=1,...,N$
9:  end while
9:  $z_{\partial v_{1}},z_{\partial v_{2}},...,z_{\partial v_{N}}$
@@
(b) M2PDE when predicting $T$.
Figure 9: Comparison of surrogate models and M2PDE + Transolver for predicting the temperature of large structures.
@@
(b) M2PDE when predicting $\varepsilon_{x}$.
Figure 10: Comparison of surrogate models and M2PDE + Transolver for predicting the strain of large structures.
(a) M2PDE when predicting $T$.
(b) M2PDE when predicting $\varepsilon_{x}$.
Figure 11: The results of M2PDE + FNO for predicting large structure.
@@ F.3 Selection of Hyperparameter of $\lambda$
This section compares the computational efficiency of M2PDE, surrogate model, and numerical programs. The time unit for each experiment is defined as the time required for a single neural network inference. These three experiments all use the most accurate model. Since the surrogate model and M2PDE both use the same network architecture and have consistent network parameters, it is assumed that the time for a single inference using these two methods is equal. The numerical programs are run on the CPU and have all been optimized to the best parallel count.
@@
The results are presented in Table 20. In experiment 1, the problem is relatively simple, and the numerical algorithm achieves efficient solutions through explicit time stepping, while the introduction of M2PDE actually reduces efficiency. However, in experiment 2, which addresses more complex problems, M2PDE achieves a 29-fold acceleration compared to numerical programs. In experiment 3, comparing the results of 16 components with 64 components, it is observed that as the computational scale increases, the acceleration effect of M2PDE becomes increasingly significant. Furthermore, when dealing with multi-component problems, the surrogate model requires iteration to ensure the convergence of solutions across all components. Due to the large number of components, the number of iterations needed significantly increases compared to multiphysics problems, resulting in higher efficiency for M2PDE. In addition, we have only compared the efficiency of single computations for all experiments. When dealing with multiple problems simultaneously, the acceleration provided by M2PDE will be even more pronounced due to the parallel nature of GPU computing.
In general, the more complex the problem, the more pronounced the acceleration effect of M2PDE becomes. In fact, the problems in experiment 2 and experiment 3 have been simplified to a certain extent, and the actual situations are even more complex. Therefore, M2PDE holds significant value in solving real-world complex engineering problems.
@@
Experiment Unit (s) Numerical program Surrogate model M2PDE Speedup
@@
In this section, we discuss the application scenarios of M2PDE for multi-component simulation from both theoretical and practical perspectives.
From a theoretical perspective, in the derivation of Section 3.2, we make an assumption: the solution on a multi-component structure is an undirected graph that satisfies the local Markov property, meaning that any two non-adjacent variables are conditionally independent given all other variables. Using this property, we derived Eq. 14. We believe this assumption is applicable to most problems because physical fields are continuous in space, and the information exchange between any two points must be transmitted through the points in between. However, there is a class of problems to which current methods cannot be directly applied, which is the PDE that requires determining eigenvalues: